• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 00:12:28 on April20,2004

    Entry number 118628

    keyword=swing shift summary


    16:00 Shift starts with no beam, recovering from "CD0 festivies"

    17:10 Beam returns. BPM 4a (1.9,0.2), 4b(-1.9,-1.5)
    FFB on.
    started run 2558. Nominally "junk", as beam just came on, but
    actually beam looks pretty good (couple of trips, current changing).

    17:25 90 uA, production starts.

    18:30 ep power supplies turned on, waiting to stabalize.

    18:42 moved to 10cm optics target. Started Doug's optics program, changed
    prescales T1=4, T3=1, beam about 5 uA.
    Z-E. worries about deadtime, figured we should prescale more. Started
    run 2562, with T1=8, T3=2. roughly 16% dead.

    19:00 moved to 24cm optics target

    19:05 Arne called: he will start energy spread measurement and OTR/SLI cross-calibration
    expected: 1 hour. Turned off ep detectors to avoid damage. Beamline HV in ep system
    tripped, so those channels were turned off as well.

    20:05 Arne called: energy spread is bad, expert on the way. Until expert arrives, we
    returned to the optics runs.

    22:05 Energy spread appears to be good, measured by OTR to be <3e-3. This is true only
    at high current, so no harp scan was done to verify setpoints.

    22:10 Jay Benesch trying some new quad values for Hall A to try to reduce trip
    rate.

    22:25 MCC called. They're ready to try the new quad settings. They remeasured
    OTR width with new setpoints to verify energy width was unchanged: ELog entry.

    22:45 Lots of trips, no win. MCC is trying to find the upper limit for current.
    Jaideep points out that we have pulsed beam... at 33 milliHertz.

    23:20 Trip rate still high...